An 80-lb weight is placed 9 ft from a fulcrum. What weight should you put 6 ft from the fulcrum to balance the lever?
Answer by checkley77(12844)

6/9=80/X
6X=9*80
6X=720
X=720/6
X=120 LB. ANS.